OpenStudy (anonymous):

Calculate the electric field at distance of 320cm from an infinitely long wire carrying a charge per unit length equal to 8.3 C/m.

OpenStudy (nali):

Use the equation E = ( q/l ) / 2 pi epsilon 0 (e0) r^2 So just plug in the q/ l

OpenStudy (nali):

So plug in q/l as 8.3 C/m since it is given as charge over length Then just use the given radius and plug it into the equation to get the electric field

OpenStudy (nali):

this the equation just in case you do not get it from my first comment \[\large E=(q/l)/ 2 \pi \epsilon 0 r^2\]